# Revisiting the question of what instantaneous normal modes tell us about liquid dynamics

### v1

@@ II.2 Simulation details
For the four monatomic systems, a FCC lattice structure of 4000 atoms with lattice constant of $a_{A}$r=$5.268\,\textup{\AA}$ Henshaw (1958), $a_{X}$e= $6.347\,\textup{\AA}$ Granfors et al. (1981), $a_{G}$a= $4.4\,\textup{\AA}$ Belashchenko (2012) and $a_{P}$b= $4.95\,\textup{\AA}$ Wang et al. (2018) was respectively used as an initial input structure for the MD simulations. For N2 and CS2 systems, the MD simulations were performed with 500 N2 and 3375 CS2 molecules in either one cubic box of dimension 30.0 Å and 80.0 Å respectively. All samples were simulated with periodic boundary conditions (PBC). In the process of samples preparation, after energy minimization was performed using steepest descent algorithm to bring the initial configuration in minimum potential, all samples relaxed at least $5$ ns (MD time step is $1.0$ fs) to reach their respective equilibrium at a series of temperatures and specific pressures based on phase diagram to keep them liquid. In this process, the isobaric and isothermal (NPT) ensemble was used with the sample size being adjusted to give target pressure. After that, the canonical (NVT) ensemble MD was conducted at different temperatures for data collections and analysis with an interval of $100$ fs during a $1$ ns simulation. The MD simulation systems of Ar and CS2 are shown in Figure 1.
Figure 1: MD simulation systems for liquid Ar and CS2. The snapshot of liquid Ar (a) at $T=100$ K, $P=40$ bar and liquid CS2 (b) at $T=300$ K, $P=1.01$ bar. Figures made with OVITO software Stukowski (2009).
@@ II.3 Definition of the physical observables
The velocity auto-correlation function (VACF), $Z(t)$, is an essential time-correlation function in physics and chemistry that provides extensive information about molecular-structural and hydrodynamic properties of matter Hansen and McDonald (2013). The Fourier transform of $Z(t)$ is defined as the density of state function, $S(\omega)$. More precisely,
@@
On the other hand, the instantaneous normal mode analysis provides an instantaneous picture of liquid dynamics that has been widely used in the past Keyes (1997a). The INM method involves calculating the normal modes of a system at a given instantaneous configuration, providing insight into the vibrational properties and stability of the system at specific moments in time. It starts from the dynamical matrix that characterizes the curvatures of the local potential energy surface. The dynamical matrix is a $3N\times 3N$ matrix whose elements are defined as follows:
@@ III Argon as a benchmark case
We take the case of Argon (Ar) as a paradigmatic simple example to study liquid properties within the INM framework. The phase diagram of Argon in the regime of interest is shown in Fig.2. The black line indicates the line of first-order phase transitions separating the liquid (light blue) and the gas (light green) phases. The critical point, indicated with a red circle, is located at $T_{c}\approx 151$ K and $P_{c}\approx 48.55$ bar Mccain and Ziegler (1967). The data for the phase diagram are collected from the National Institute of Standards and Technology (NIST) database Standards (2005). The pink region is the supercritical phase beyond the critical point. The horizontal red isobar lines indicate some of the regions that have been numerically investigated in this work. Other investigated isobar lines with much higher pressure are not displayed in Fig.2. Finally, the blue circles indicate the onset of the dynamical crossover (Frenkel line) that will be discussed in detail below.
@@
We start our analysis by considering the behavior of the velocity auto-correlation function (VACF) $Z(t)$ and the pair distribution function (PDF) $g(r)$ that are shown in panels a and b of Fig.3 respectively. At high temperature, deep in the gas regime, $Z(t)$ decays monotonically and its form is well approximated by an exponential function $\exp(-\gamma t)$ (shown with black dashed line), as predicted by kinetic theory under the assumption of pure Brownian motion Hansen and McDonald (2013). Upon decreasing the temperature, $Z(t)$ remains a monotonically decreasing function but it is not anymore well approximated by a pure exponential function. This indicates the onset of correlations and the tendency towards dense liquid dynamics. A big jump is observed between $140$ K and $130$ K that corresponds approximately to the boiling temperature of Argon at this pressure. Even in the liquid phase, at least up to a certain “critical” temperature, the velocity auto-correlation function remains a monotonically decreasing function.
More importantly, as emphasized in the inset of panel a in Fig.3, the velocity-autocorrelation function displays a drastic change of behavior around $115$ K where it starts developing a minimum that transforms into marked oscillations by decreasing the temperature further. This drastic change of behavior corresponds to a transition between gas-like dynamics to liquid-like dynamics and it is known in the literature as the Frenkel line Brazhkin and Trachenko (2012); Brazhkin et al. (2012, 2013). The Frenkel line has been studied in several liquids in the supercritical phase Yoon et al. (2018); Bolmatov et al. (2013a, 2014); Prescher et al. (2017); Bolmatov et al. (2015); Fomin et al. (2018); Proctor et al. (2019); Fomin et al. (2015); Brazhkin et al. (2011); Cockrell et al. (2021), but also in complex dusty plasmas Huang et al. (2023); Ge et al. (2023) and active granular fluids Jiang et al. (2024). Here, we report the existence of this dynamical crossover also below the critical point, away from the supercritical region. This confirms, as already noted in the previous literature, that the Frenkel line does not emanate from the critical point and it persists also in the liquid phase. We have tracked the point at which this dynamical crossover appears for three different pressures. The results are indicated with blue symbols in the phase diagram, Fig.2. From there, it is clear that the Frenkel line extends into the liquid phase and goes almost parallel to the line of first-order phase transitions below the critical point.
@@
In order to investigate structural properties, in panel b of Fig.3 we have computed the pair distribution function $g(r)$ in the same range of temperatures and constant pressure $P=40$ bar. At low temperatures, we observe the typical behavior of a dense liquid with a first strong peak around $4\,\textup{\AA}$ and other weaker peaks around $7\,\textup{\AA}$, $10.5\,\textup{\AA}$ and so on. By increasing the distance $r$, the peaks disappear signaling the absence of long-range order, as one might expect in a liquid. As shown in the inset of panel b of Fig.3, by increasing the temperature but still below the boiling temperature of $\approx 140$ K, we observe that the higher peaks (above the first two) corresponding to intermediate-range order disappear. For example, as indicated by the black arrows, the intensity of the third peak located at approximately $r\approx 10.5\,\textup{\AA}$ vanishes around $115-120$ K. Interestingly, these temperatures are extremely close to the location at which oscillations in the VACF $Z(t)$ disappear (panel a in Fig.3). As already proposed in Bolmatov et al. (2013b), this suggests the existence of a structural change in the intermediate-range order that might be associated to the Frenkel line.
@@
At approximately $95$ K, the system enters the liquid phase and the behavior of the density of state function $S(\omega)$ changes drastically. First, the zero frequency value of $S(\omega)$ is not anymore zero as a sign of a finite self-diffusion constant that grows with temperature. Moreover, $S(\omega)$ is not anymore monotonic. It first decays monotonically up to a minimum located around $\omega=1$ ps-1, then, it grows up to a peak at approximately $2.5$ ps-1 and it then decays monotonically again. This behavior is typical of dense liquids, see for example Fig.1 in Ref. Lin et al. (2003). By increasing further the temperature, both the minimum and the peak at low frequency disappear and the density of state function becomes a monotonically decaying function as expected for dilute liquids and gases. By increasing temperature, $S(\omega)$ gets closer to the Lorentzian function predicted by kinetic theory, a sign of the complete lack of correlations.
We can already anticipate some clear differences and similarities between the stable INM density of states, $\rho(\omega)$ with $\omega>0$, and the density of state function $S(\omega)$. For brevity, we will indicate the positive ($\omega>0$) branch of $\rho(\omega)$ as $\rho_{s}(\omega)$ in the rest of the manuscript, where the subscript stands for “stable”. Let us start from the solid phase near melting (black lines). Even there, $\rho_{s}(\omega)$ and $S(\omega)$ are profoundly different below $\approx 4$ ps-1. In particular, while $\rho_{s}(\omega)$ is linear in frequency, $S(\omega)$ is quadratic, as Debye’s model predicts for solids. The origin of this difference can be found in the existence of a large number of unstable modes, that also persists in the heated crystal away from the zero temperature harmonic limit. These remaining unstable modes possibly correspond to local anharmonicities that remain present in crystals at finite temperature. Above a certain critical frequency, $\rho_{s}(\omega)$ and $S(\omega)$ become more similar, showing the same qualitative trend. We notice that the positions of the peak in $\rho_{s}(\omega)$ and $S(\omega)$ are compatible in the solid phase. We also emphasize that the fact that the INM analysis does not capture correctly the long-time dynamics is not a surprise and it is a direct consequence of the instantaneous approximation adopted. The differences between $\rho_{s}(\omega)$ and $S(\omega)$ are expected to vanish by decreasing the temperature or increasing the pressure, as discussed in detail in Seeley et al. (1992). Below, we will show this trend concretely for another liquid system, Gallium.
@@
Figure 5: (a) The fraction of unstable modes $f_{u}(T)$ and the self-diffusion constant $D(T)$ as a function of temperature for Argon at $P=40$ bar. The melting temperature is between $92$ K and $95$ K (dashed red line) and the Frenkel line is located approximately around $115$ K (dashed blue line). (b) The slope of the low-frequency linear scaling of the INM DOS $a(T)$ for the same system and conditions.
@@
We now move to discuss in detail the features of the INM density of states. We are mainly interested in the fraction of unstable modes $f_{u}(T)$ defined in Eq.(9) and the coefficient of the low-frequency linear scaling $a(T)$, as determined by fitting the INM DOS using Eq.(10). The fraction of unstable modes is shown as a function of temperature in panel a of Fig.5. $f_{u}$ grows with temperature, as one might expect for a parameter that quantifies the “fluidity” of a system. As already anticipated, $f_{u}$ is not zero in the solid phase near melting but accounts for approximately $10\%$ of the INM. $f_{u}$ displays a sharp peak at the melting temperature between $92$ K and $95$ K, where it changes abruptly from $\approx 12\%$ to $\approx 28\%$. This type of behavior was already observed in the past literature Gezelter et al. (1997). Above melting, in the liquid phase, the fraction of unstable modes grows with temperature following an approximately linear trend and it does not show any feature at the dynamical crossover (Frenkel line) reported in panel a of Fig.3. In panel a of Fig.5, we show the self-diffusion constant $D$ as well. Below the melting transition (dashed red line), but not too far from it, the self-diffusion constant becomes very small, as one might expect in crystals Mehrer (2007). This implies that most of the corresponding non-vanishing unstable modes should not be of diffusive type. In the past, these remaining unstable modes have been associated to so-called “false barriers” Gezelter et al. (1997) or simply described as localized unstable modes Bembenek and Laird (1995); Clapa et al. (2012b). Above the melting temperature, the diffusion constant becomes large and grows with temperature monotonically. Within the precision of our data, we do not observe any clear change of behavior across the Frenkel temperature around $115$ K (blue dashed line), despite $D$ seems to display a quite sharp jump around that temperature. A more in detail investigation is needed to further clarify this point.
@@
In panel b of Fig.5, we show the temperature dependence of the linear scaling of the INM DOS, $a(T)$. This scaling coefficient is the same for the stable and unstable parts of the INM spectrum. $a(T)$ grows monotonically with temperature and displays a sharp peak around the melting temperature. Interestingly, its trend between $95$ K and $100$ K seems to follow the behavior of the self-diffusion constant. In particular, in that range of temperatures both $D$ and $a(T)$ are approximately constant before approaching a more steady growth. On the other hand, we do not observe any feature in $a(T)$ upon crossing the Frenkel crossover.
@@
Figure 6: (a) The momentum gap ($k$-gap) of collective shear waves $k_{g}$ as a function of temperature in the subcritical liquid phase of Argon between $95$ K and $115$ K. The data are taken from Ref. Yang et al. (2017). For convenience, the data of the fraction of unstable INM $f_{u}$ are reproduced in the same panel. (b) The $k$-gap as a function of the fraction of unstable modes $f_{u}$ and as a function of the self-diffusion constant $D$.
@@
The presence of unstable modes, $f_{u}\neq 0$, has been often identified as a hallmark of fluidity and a typical feature of liquids. As we have already observed, this is not that simple since a portion of unstable modes survive in the solid phase (at least near to melting) and a more refined classification of the unstable INM is needed to connect them directly with “fluidity” (see e.g. Bembenek and Laird (1995); Clapa et al. (2012a); Sciortino and Tartaglia (1997); La Nave et al. (2000, 2001); Chowdhary and Keyes (2002); Gezelter et al. (1997); Li and Keyes (1997) for the several criteria proposed). Nevertheless, we find it interesting to compare the fraction of unstable modes and the self-diffusion constant to another measure of “fluidity” known as the $k$-gap. As originally advocated by Maxwell and Frenkel Frenkel (1946), and later extensively re-discussed Trachenko and Brazhkin (2015b); Baggioli et al. (2020), collective shear waves in liquids do not display a linear dispersion relation, as in solids, due to the vanishing of the static shear modulus. On the other hand, the real part of their dispersion (i.e., their energy) follows an approximate square-root form $\mathrm{Re}(\omega)=v\sqrt{k^{2}-k_{g}^{2}}$ where $k_{g}$ is known in the literature as $k$-gap. This structure has been confirmed in several simulations (e.g., Yang et al. (2017); Khusnutdinoff et al. (2020); Ge et al. (2023)) and has been also observed in few experimental systems Nosenko et al. (2006); Jiang et al. (2024). Interestingly, the value of $k_{g}$ is rather universal when compared to the Wigner-Seitz radius Baggioli (2021). $k_{g}$ is expected to vanish at the melting transition, below which the collective shear waves reacquire their propagating nature as predicted by hydrodynamics Martin et al. (1972) and elasticity theory Chaikin and Lubensky (1995). This suggests that somehow $k_{g}$ should relate to the self-diffusion constant and the fraction of “diffusive” unstable INM that also vanish in the solid phase. Nevertheless, to the best of our knowledge, how this net of relations is concretely realized has never been investigated.
@@
Taking advantage of the existing data for $k_{g}$ for subcritical liquid Ar Yang et al. (2017), we have attempted a raw comparison among these quantities. The results are shown in Fig.6. In panel a, we have presented the data for $k_{g}$ as a function of temperature in the liquid phase. As expected, $k_{g}$ grows monotonically with temperature compatible with the idea that $k_{g}\propto 1/\tau_{m}$, with $\tau_{m}$ being the Maxwell time. The latter decreases with temperature and make $k_{g}$ to increase with $T$. For comparison, in the same panel, we also show the fraction of unstable modes as a function of temperature, to highlight the very similar trend. In panel b of Fig. 6, we plot $k_{g}$ as a function of $D$ and $f_{u}$. In both cases, we observed a very strong linear correlation between these quantities. This is consistent with the idea that all of them are a measure of “fluidity”.

### v2

@@ II.2 Simulation details
For the four monatomic systems, a Face-Centered Cubic (FCC) lattice structure of 4000 atoms with lattice constant of $a_{A}$r=$5.268\,\textup{\AA}$ [77], $a_{X}$e= $6.347\,\textup{\AA}$ [78], $a_{G}$a= $4.4\,\textup{\AA}$ [67] and $a_{P}$b= $4.95\,\textup{\AA}$ [68] was respectively used as an initial input structure for the MD simulations. For N2 and CS2 systems, the MD simulations were performed with 500 N2 and 3375 CS2 molecules in either one cubic box of dimension 30.0 Å and 80.0 Å respectively. All samples were simulated with periodic boundary conditions (PBC). In the process of samples preparation, after energy minimization was performed using steepest descent algorithm to bring the initial configuration in minimum potential, all samples relaxed at least $5$ ns (MD time step is $1.0$ fs) to reach their respective equilibrium at a series of temperatures and specific pressures based on phase diagram to keep them liquid. In this process, the isobaric and isothermal (NPT) ensemble was used with the sample size being adjusted to give target pressure. After that, the canonical (NVT) ensemble MD was conducted at different temperatures for data collections and analysis with an interval of $100$ fs during a $1$ ns simulation. The MD simulation systems of Ar and CS2 are shown in Figure 1.
Figure 1: MD simulation systems for liquid Ar and CS2. The snapshot of liquid Ar (a) at $T=100$ K, $P=40$ bar and liquid CS2 (b) at $T=300$ K, $P=1.01$ bar. Figures made with OVITO software [79].
The pressures of all of our simulated systems are provided in the Table 3. In our simulation systems, the temperature range for argon is the most extensive, covering all three phases: solid, liquid, and gas. The approximate melting and boiling points for argon are provided in Table 3. For xenon, plumbum, and gallium, the simulated temperature range includes both solid and liquid phases, with the melting temperature presented in Table 3 as well. It is important to highlight that, since the phase transition dynamics are not the focus of our work, our simulations were performed at fixed temperature rather than simulating the phase transition process itself. The melting point was determined based on the system’s state at different temperatures rather than through direct observation of the transition dynamics (see details in Appendix .1). As a result, the identified melting points are not accurate and may slightly differ from the experimental measurements. To provide context, experimental data from the literature are also included in the table for comparison. For N2 and CS2 systems, we focused on simulating the liquid phase; both the melting and boiling points for these two systems presented in Table 3 are taken from experimental data.
Table 3: Melting and boiling temperatures for the systems considered in this work.
System Pressure (bar) Melting temperature (K) Boiling temperature (K)
Ar 40 92 130
Xe 1.01 145, 161 [80] 165 [80]
Pb 1.01 740, 600 [81] 2013 [81]
Ga 1.01 320, 303 [82] 2676 [83]
N2 70.92 63 [84] 127 [82]
CS2 1.01 161 [82] 319 [82]
@@ II.3 Definition of the physical observables
The velocity auto-correlation function (VACF), $Z(t)$, is an essential time-correlation function in physics and chemistry that provides extensive information about molecular-structural and hydrodynamic properties of matter [85]. The Fourier transform of $Z(t)$ is defined as the density of state function, $S(\omega)$. More precisely,
@@
On the other hand, the instantaneous normal mode analysis provides an instantaneous picture of liquid dynamics that has been widely used in the past [12]. The INM method involves calculating the normal modes of a system at a given instantaneous configuration, providing insight into the vibrational properties and stability of the system at specific moments in time. It starts from the dynamical matrix that characterizes the curvatures of the local potential energy surface. The dynamical matrix is a $3N\times 3N$ matrix whose elements are defined as follows:
@@ III Argon as a benchmark case
We take the case of Argon (Ar) as a paradigmatic simple example to study liquid properties within the INM framework. The phase diagram of Argon in the regime of interest is shown in Fig.2. The black line indicates the line of first-order phase transitions separating the liquid (light blue) and the gas (light green) phases. The critical point, indicated with a red circle, is located at $T_{c}\approx 151$ K and $P_{c}\approx 48.55$ bar [86]. The data for the phase diagram are collected from the National Institute of Standards and Technology (NIST) database [82]. The pink region is the supercritical phase beyond the critical point. The horizontal red isobar lines indicate some of the regions that have been numerically investigated in this work. Other investigated isobar lines with much higher pressure are not displayed in Fig.2.
@@
We start our analysis by considering the behavior of the velocity auto-correlation function (VACF) $Z(t)$ and the pair distribution function (PDF) $g(r)$ that are shown in panels a and b of Fig.3 respectively. At high temperature, deep in the gas regime, $Z(t)$ decays monotonically and its form is well approximated by an exponential function $\exp(-\gamma t)$ (shown with black dashed line), as predicted by kinetic theory under the assumption of pure Brownian motion [85]. Upon decreasing the temperature, $Z(t)$ remains a monotonically decreasing function but it is not anymore well approximated by a pure exponential function. This indicates the onset of correlations and the tendency towards dense liquid dynamics. A big jump is observed between $140$ K and $130$ K that corresponds approximately to the boiling temperature of Argon at this pressure. Even in the liquid phase, at least up to a certain “critical” temperature, the velocity auto-correlation function remains a monotonically decreasing function.
More importantly, as emphasized in the inset of panel a in Fig.3, the velocity-autocorrelation function displays a drastic change of behavior around $115$ K where it starts developing a minimum that transforms into marked oscillations by decreasing the temperature further. This drastic change of behavior corresponds to a transition between gas-like dynamics to liquid-like dynamics and it is known in the literature as the Frenkel line [87, 88, 89]. The Frenkel line has been studied in several liquids in the supercritical phase [90, 91, 92, 93, 94, 95, 96, 97, 98, 99], but also in complex dusty plasmas [100, 101] and active granular fluids [102].
@@
In order to investigate structural properties, in panel b of Fig.3 we have computed the pair distribution function $g(r)$ in the same range of temperatures and constant pressure $P=40$ bar. At low temperatures, we observe the typical behavior of a dense liquid with a first strong peak around $4\,\textup{\AA}$ and other weaker peaks around $7\,\textup{\AA}$, $10.5\,\textup{\AA}$ and so on. By increasing the distance $r$, the peaks disappear signaling the absence of long-range order, as one might expect in a liquid. As shown in the inset of panel b of Fig.3, by increasing the temperature but still below the boiling temperature of $\approx 140$ K, we observe that the higher peaks (above the first two) corresponding to intermediate-range order disappear. For example, as indicated by the black arrows, the intensity of the third peak located at approximately $r\approx 10.5\,\textup{\AA}$ vanishes around $115-120$ K. Interestingly, these temperatures are extremely close to the location at which oscillations in the VACF $Z(t)$ disappear (panel a in Fig.3). As already proposed in [103], this suggests the existence of a structural change in the intermediate-range order that might be associated to the Frenkel line.
@@
At approximately $95$ K, the system enters the liquid phase and the behavior of the density of state function $S(\omega)$ changes drastically. First, the zero frequency value of $S(\omega)$ is not anymore zero as a sign of a finite self-diffusion constant that grows with temperature. Moreover, $S(\omega)$ is not anymore monotonic. It first decays monotonically up to a minimum located around $\omega=1$ ps-1, then, it grows up to a peak at approximately $2.5$ ps-1 and it then decays monotonically again. This behavior is typical of dense liquids, see for example Fig.1 in Ref. [60]. By increasing further the temperature, both the minimum and the peak at low frequency disappear and the density of state function becomes a monotonically decaying function as expected for dilute liquids and gases. By increasing temperature, $S(\omega)$ gets closer to the Lorentzian function predicted by kinetic theory, a sign of the complete lack of correlations.
We can already anticipate some clear differences and similarities between the stable INM density of states, $\rho(\omega)$ with $\omega>0$, and the density of state function $S(\omega)$. For brevity, we will indicate the positive ($\omega>0$) branch of $\rho(\omega)$ as $\rho_{s}(\omega)$ in the rest of the manuscript, where the subscript stands for “stable”. Let us start from the solid phase near melting (black lines). Even there, $\rho_{s}(\omega)$ and $S(\omega)$ are profoundly different below $\approx 4$ ps-1. In particular, while $\rho_{s}(\omega)$ is linear in frequency, $S(\omega)$ is quadratic, as Debye’s model predicts for solids. The origin of this difference can be found in the existence of a large number of unstable modes, that also persists in the heated crystal away from the zero temperature harmonic limit. These remaining unstable modes possibly correspond to local anharmonicities that remain present in crystals at finite temperature. Above a certain critical frequency, $\rho_{s}(\omega)$ and $S(\omega)$ become more similar, showing the same qualitative trend. We notice that the positions of the peak in $\rho_{s}(\omega)$ and $S(\omega)$ are compatible in the solid phase. We also emphasize that the fact that the INM analysis does not capture correctly the long-time dynamics is not a surprise and it is a direct consequence of the instantaneous approximation adopted. The differences between $\rho_{s}(\omega)$ and $S(\omega)$ are expected to vanish by decreasing the temperature or increasing the pressure, as discussed in detail in [16]. Below, we will show this trend concretely for another liquid system, Gallium.
@@
Figure 5: (a) The fraction of unstable modes $f_{u}(T)$ and the self-diffusion constant $D(T)$ as a function of temperature for Argon at $P=40$ bar. The melting temperature is between $92$ K and $95$ K (dashed red line). (b) The slope of the low-frequency linear scaling of the INM DOS $a(T)$ for the same system and conditions.
@@
We now move to discuss in detail the features of the INM density of states. We are mainly interested in the fraction of unstable modes $f_{u}(T)$ defined in Eq.(9) and the coefficient of the low-frequency linear scaling $a(T)$, as determined by fitting the INM DOS using Eq.(10). The fraction of unstable modes is shown as a function of temperature in panel a of Fig.5. $f_{u}$ grows with temperature, as one might expect for a parameter that quantifies the “fluidity” of a system. As already anticipated, $f_{u}$ is not zero in the solid phase near melting but accounts for approximately $10\%$ of the INM. $f_{u}$ displays a sharp peak at the melting temperature between $92$ K and $95$ K, where it changes abruptly from $\approx 12\%$ to $\approx 28\%$. This type of behavior was already observed in the past literature [21]. Above melting, in the liquid phase, the fraction of unstable modes grows with temperature following an approximately linear trend. As mentioned above in the section describing the simulation details, it is important to clarify that we did not simulated directly the dynamical melting process. Rather, we have located the melting point by looking at changes in the properties of the system at fixed temperature (see Appendix .1 for more details). Our analysis suggests that the system is in a solid state below $T=92$ K but transitions to a liquid state at $95$ K, signifying the occurrence of a phase transition within that range of temperatures.
In panel a of Fig.5, we show the self-diffusion constant $D$ as well. Below the melting transition (dashed red line), but not too far from it, the self-diffusion constant becomes very small, as one might expect in crystals [104]. This implies that most of the corresponding non-vanishing unstable modes should not be of diffusive type. In the past, these remaining unstable modes have been associated to so-called “false barriers” [21] or simply described as localized unstable modes [63, 25]. Above the melting temperature, the diffusion constant becomes large and grows with temperature monotonically.
@@
In panel b of Fig.5, we show the temperature dependence of the linear scaling of the INM DOS, $a(T)$. This scaling coefficient is the same for the stable and unstable parts of the INM spectrum. $a(T)$ grows monotonically with temperature and displays a sharp peak around the melting temperature. Interestingly, its trend between $95$ K and $100$ K seems to follow the behavior of the self-diffusion constant. In particular, in that range of temperatures both $D$ and $a(T)$ are approximately constant before approaching a more steady growth.
@@
Figure 6: (a) The momentum gap ($k$-gap) of collective shear waves $k_{g}$ as a function of temperature in the subcritical liquid phase of Argon between $95$ K and $115$ K. The data are taken from Ref. [105]. For convenience, the data of the fraction of unstable INM $f_{u}$ are reproduced in the same panel. (b) The $k$-gap as a function of the fraction of unstable modes $f_{u}$ and as a function of the self-diffusion constant $D$.
@@
The presence of unstable modes, $f_{u}\neq 0$, has been often identified as a hallmark of fluidity and a typical feature of liquids. As we have already observed, this is not that simple since a portion of unstable modes survive in the solid phase (at least near to melting) and a more refined classification of the unstable INM is needed to connect them directly with “fluidity” (see e.g. [63, 25, 106, 107, 108, 23, 21, 42] for the several criteria proposed). Nevertheless, we find it interesting to compare the fraction of unstable modes and the self-diffusion constant to another measure of “fluidity” known as the $k$-gap. As originally advocated by Maxwell and Frenkel [109] (see also fundamental contributions by Zwanzig [49]), and later extensively re-discussed [56, 110], collective shear waves in liquids do not display a linear dispersion relation, as in solids, due to the vanishing of the static shear modulus. On the other hand, the real part of their dispersion (i.e., their energy) follows an approximate square-root form $\mathrm{Re}(\omega)=v\sqrt{k^{2}-k_{g}^{2}}$ where $k_{g}$ is known in the literature as $k$-gap. This structure has been confirmed in several simulations (e.g., [105, 111, 101]) and has been also observed in few experimental systems [112, 102].
Interestingly, the $k$-gap dispersion relation of shear waves in liquids might also provide an explanation for their linear in frequency DOS at low frequency. Indeed, as explicitly derived in [113, 114], a $k$-gap $k_{g}=1/(2v\tau_{m})$ (with $\tau_{m}$ the Maxwell relaxation time) implies a DOS of the form
$g(\omega)\propto\frac{\omega^{2}}{v^{3}}\sqrt{1+\frac{1}{4\omega^{2}\tau_{m}^{2}}}.$ (13)
At low frequencies, i.e. $\omega\ll 1/(2\tau_{m})$, the DOS exhibits a linear scaling $g(\omega)\propto\omega/\tau_{m}$ that is compatible with the findings from the INM analysis. We notice that in that low-frequency regime shear fluctuations are not propagating but rather overdamped (i.e., having imaginary frequency or complex frequency with imaginary part larger than the real). This suggests a connection between linear in frequency DOS, unstable modes and $k$-gap that deserves further attention. We also notice that, despite the INM analysis correctly captures the linear scaling of the DOS at low-frequencies (as predicted as well by $k$-gap theory, Eq. 13), it overestimates the importance of unstable modes. This is specially evident close to the melting temperature where the dispersion of shear waves is almost linear $\omega\approx vk$ (see for example [15, 115, 116, 117, 118]) and therefore the linear contribution is expected to be very small. At this stage, the origin of this problem is not evident but it is likely related to the inherent assumptions of the INM analysis (instantaneous and harmonic).
We reiterate again that the linear scaling of the liquid DOS at low frequency comes from overdamped non-propagating modes. This interpretation is in nice agreement with the above computation from the $k$-gap dispersion, showing that the modes contributing to that linear regime have $\omega\ll 1/\tau_{m}$ and, hence, are not propagating wave-like excitations.
As already mentioned, $k_{g}$ is expected to vanish at the melting transition, below which the collective shear waves reacquire their propagating nature as predicted by hydrodynamics [119] and elasticity theory [120]. This suggests that somehow $k_{g}$ should relate to the self-diffusion constant and the fraction of “diffusive” unstable INM that also vanish in the solid phase. Nevertheless, to the best of our knowledge, how this net of relations is concretely realized has never been investigated.
@@
Taking advantage of the existing data for $k_{g}$ for subcritical liquid Ar [105], we have attempted a raw comparison among these quantities. The results are shown in Fig.6. In panel a, we have presented the data for $k_{g}$ as a function of temperature in the liquid phase. As expected, $k_{g}$ grows monotonically with temperature compatible with the idea that $k_{g}\propto 1/\tau_{m}$, with $\tau_{m}$ being the Maxwell time. The latter decreases with temperature and make $k_{g}$ to increase with $T$. For comparison, in the same panel, we also show the fraction of unstable modes as a function of temperature, to highlight the very similar trend. In panel b of Fig. 6, we plot $k_{g}$ as a function of $D$ and $f_{u}$. In both cases, we observed a very strong linear correlation between these quantities. This is consistent with the idea that all of them are a measure of “fluidity”.
@@
Before concluding, we notice that an approximately linear relation between the diffusion constant $D$ and the $k$-gap $k_{g}$ could be explained by combining $k$-gap theory with the Stokes-Einstein relation. In fact, the latter implies $D\propto T/\eta=T/(G_{\infty}\tau_{m})$. Using $k_{g}=1/(2v\tau_{m})$, one immediately gets $D\propto k_{g}$, that is compatible with the trend in the simulation data, Fig. 6b.